You are on a planet with an unknown acceleration due to gravity. You drop an object from rest and a height 1.1 metres, and measure the final speed 5.6 m/s as the object hits the ground. Using the Principle of Conservation of Energy, calculate the magnitude of the acceleration due to gravity. Use 2 sf in your answer.
Answer #1

Solution) height , h = 1.1 m

Speed ,V = 5.6 m/s

Acceleration due to gravity , a = ?

Initially when object is thrown from certain height it possess potential energy(PE) and when it falls it possess kinetic energy(KE) .

So applying principle of conservation of energy

PE = KE

M(a)h = (1/2)(M)(V^2)

a = (V^2)/(2h)

a = (5.6^2)/(2×1.1)

a = 14.25 m/s^2

Upto two significant figures

So acceleration a = 14 m/s^2